OpenStudy (anonymous):

x + 5y = -2 2x + y = 5 The point of intersection of the lines has a y-coordinate of _____.

OpenStudy (anonymous):

Welcome 2 openstudy

OpenStudy (anonymous):

x + 5y = -2 -->(-2) 2x + y = 5 ------------- -2x - 10y = 4 (result of multiplying by -2) 2x + y = 5 -------------add -9y = 9 y = -1 your y coordinate at the point of intersection is (0,-1)

OpenStudy (anonymous):

your y coordinate is -1

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

can you help me with another question?